Name: _01. Which of the choices below constitutes a simultaneous solution to these equations? (3pts.) (1) 3X + 2Y = 10 and (2 need all questions answered as soon as possible i have an operations management test in the morning and really need these to study. im having a lot of trouble answering the problems and need to see hpw it is done. thank you.
Answer #1

1) Ans A

X = 6, Y = -4

3X + 2Y = 3*6 + 2*-4 = 10

5X + 4Y = 5*6 + 4*-4 = 14

So, both equations are satisfied